# On Prior Distributions for Orthogonal Function Sequences [^1] August 21, 2025

### v1

@@ Introduction
$f_{1}(t)$ $={\rm I}\Big{(}0\leq t<\frac{1}{2}\Big{)}-{\rm I}\Big{(}\frac{1}{2}\leq t\leq 1\Big{)},\ \ \ \ f_{2}(t)=\sqrt{2}\left\{{\rm I}\Big{(}0\leq t<\frac{1}{4}\Big{)}-{\rm I}\Big{(}\frac{1}{4}\leq t<\frac{1}{2}\Big{)}\right\},$
$\hskip 76.82234ptf_{3}(t)=\sqrt{2}\left\{{\rm I}\Big{(}\frac{1}{2}\leq t<\frac{3}{4}\Big{)}-{\rm I}\Big{(}\frac{3}{4}\leq t\leq 1\Big{)}\right\}.$
@@
for the Haar wavelet scenario, where ${\rm I}(\cdot)$ is an indicator function. The latent scores for the first three components were generated independently from normal distributions with decreasing variances to reflect dominant leading components: $\xi_{1}\sim{\mathcal{N}}(0,1)$, $\xi_{2}\sim{\mathcal{N}}(0,0.7^{2})$, and $\xi_{3}\sim{\mathcal{N}}(0,0.5^{2})$. The observed functional data $Y_{i}(t)$ were then obtained by adding Gaussian noise with standard deviation $\sigma=1$ to the true signal.
@@
For the generated dataset, we applied the proposed adaptive orthogonal prior with a global inverse-gamma prior on $\tau^{2}$ (AOP-G), and local inverse-gamma priors on each $\tau_{k}^{2}$ (AOP-L), where cubic B-splines with an intercept are used for basis functions. For comparison, we adopted a non-orthogonal prior (NO) which uses a standard normal prior for each coefficients, and a non-orthogonal shrinkage prior (NO-S) applying the horseshoe prior (Carvalho et al., 2009) for coefficients of basis functions to reduce the number of irrelevant principal functions. Moreover, we employed nearly mutually orthogonal (NeMO) processes (Matuk et al., 2022). The (largest) number of possible components was set to $K=10$ in all the methods, and the number of basis functions in NO, NO-S, AOP-G and AOP-L was set to $L=12$, noting that the results were not sensitive to the choice of the different value of $L$. For NeMO, the threshold value for orthogonal constraint is set to $10^{-4}$, following the default setting of the R package provided by the author’s GitHub repository. Posterior inference was conducted via MCMC, running for 3,000 iterations after discarding 2,000 burn-in iterations.
@@
We first show the result with a single simulation data under scenario (i). In Figure 2, we visualize the estimated principal functions and heatmaps of their inner products, obtained by “NO” (non-orthogonal) and “AOP-G” (orthogonal) methods. Overall, the results demonstrated that our proposed orthogonal priors successfully enforced near-orthogonality, effectively selected the number of components, and improved reconstruction accuracy while maintaining reasonable uncertainty quantification. In particular, the orthogonal approaches accurately estimated the true number of principal components (three), with the remaining components being automatically shrunk toward zero. This shrinkage behavior is evident in the estimated principal functions, where only the leading three functions show substantial variation, while the others are effectively eliminated. Furthermore, the heatmaps of inner products provide a clear contrast between the methods: under the proposed orthogonal priors, the off-diagonal elements are nearly zero, indicating that the estimated functions are nearly orthogonal. In contrast, the non-orthogonal methods exhibit large off-diagonal inner products, implying strong correlations among components and reducing interpretability. These findings highlight the advantage of orthogonality-inducing priors in achieving interpretable, parsimonious representations in functional principal component analysis.
@@
    Number of effective components (NC): The number of components with sufficiently large variations, defined as ${\rm NC}=\sum_{k=1}^{K}{\rm I}\big{\{}\int\widehat{f}_{k}^{2}(t)dt>\epsilon\big{\}}$, where $\widehat{f}_{k}(t)\ (k=1,\ldots,K)$ is the posterior mean of the principal functions. We set $\epsilon=0.1$ in this study.
@@
Tables 1 and 2 summarize the average number of estimated principal components and the orthogonality measures, respectively, based on 200 Monte Carlo replications. Table 1 shows that the non-orthogonal model without shrinkage (NO) consistently overestimates the number of components, returning the maximum value of $K=10$ in all cases. Although the non-orthogonal model with shrinkage (NO-S) reduces the number of estimated components, it still tends to overfit, especially when the sample size is small. In contrast, the adaptive orthogonal prior approaches (AOP-G and AOP-L) provide NC values close to the true number of components, demonstrating the effectiveness of orthogonality-inducing priors in promoting sparsity. On the other hand, NeMO tends to produce NC values larger than the true number of components. Table 2 presents the orthogonality measures based on the sum of absolute off-diagonal inner products. The proposed orthogonal priors yield near-zero values, indicating that the estimated principal functions are nearly orthogonal. On the other hand, both NO and NO-S exhibit large off-diagonal values, especially in small-sample settings, reflecting strong dependencies among components. Also, it is observed that the OG values of NeMo are slightly larger than those of AOP-G and AOP-L. These results confirm that the proposed priors not only improve parsimony but also enhance interpretability by ensuring near-orthogonal decomposition.
@@
Table 1: Average values of estimated number of principal component functions (NC), based on 200 Monte Carlo replications. The standard deviations are given in the parenthesis. True number of components is 3.
Scenario $n$ NO NO-S NeMO AOP-G AOP-L
50 10.0 (0.00) 5.84 (1.63) 7.12 (1.14) 3.31 (0.56) 3.43 (0.59)
1 100 10.0 (0.00) 5.75 (1.53) 7.44 (1.05) 3.14 (0.37) 3.15 (0.39)
200 10.0 (0.07) 6.25 (1.67) 7.05 (1.08) 3.10 (0.36) 3.28 (0.51)
50 10.0 (0.00) 3.30 (0.62) 7.09 (1.17) 3.31 (0.52) 3.32 (0.53)
2 100 10.0 (0.00) 3.09 (0.32) 7.21 (1.25) 3.04 (0.18) 3.02 (0.14)
200 10.0 (0.00) 3.26 (0.48) 7.00 (1.03) 3.00 (0.00) 3.08 (0.26)
Table 2: Average values of the orthogonality measure (OG) of principal functions, based on 200 Monte Carlo replications. The standard deviations are given in the parenthesis.
Scenario $n$ NO NO-S NeMO AOP-G AOP-L
50 16.5 (3.59) 2.45 (1.06) 0.85 (0.21) 0.07 (0.14) 0.07 (0.11)
1 100 22.2 (5.05) 3.11 (1.46) 0.87 (0.19) 0.01 (0.03) 0.04 (0.24)
200 30.9 (8.15) 4.61 (3.11) 0.84 (0.19) 0.01 (0.01) 0.34 (0.67)
50 9.62 (2.22) 2.24 (1.33) 0.63 (0.22) 0.08 (0.16) 0.09 (0.14)
2 100 14.5 (3.66) 2.50 (1.56) 0.65 (0.21) 0.01 (0.04) 0.02 (0.07)
200 20.7 (5.95) 2.92 (2.04) 0.61 (0.16) 0.00 (0.00) 0.12 (0.39)
@@
We demonstrate the usefulness of the proposed orthogonal prior in Bayesian FPCA, using the population data collected by NTT Docomo Inc., which is the predominant mobile company in Japan. Following the previous studies (Wakayama et al., 2025; Wakayama and Sugasawa, 2024), we focus on the seven special wards of Tokyo metropolitan area and 500-meter mesh grids covering the seven wards. While the original data were provided as three-dimensional arrays, representing locations, time points, and dates, we use the aggregated data consisting of average hourly population counts in February 2019 on weekdays, which results in $n=452$ functional observations with $m_{i}\equiv m=24$ discrete observed points. For $i=1,\ldots,n$, let $Y_{i}^{\ast}(t)$ be the aggregated functional observation, and we use the scaled version $Y_{i}(t)$ defined as $Y_{i}(t)=Y_{i}^{\ast}(t)/\sqrt{m^{-1}\sum_{j=1}^{m}Y_{i}^{\ast}(t_{j})^{2}}$. Figure 3 shows the scaled functional data, which clearly exhibits patterns indicate strong temporal variation, which motivates the use of a low-rank functional representation via FPCA.
@@
We fitted the Bayesian FPCA with four priors, NO, NO-S, NeMO and AOP-G, to these functional data. We adopted the same settings for the prior distributions and tuning parameters as used in Section 4, such as $K=10$ (the largest number of principal functions) and $L=12$ (the number of basis functions). We generated 5,000 iterations and 5,000 burn-in. In Figure 4, we present the posterior distribution of $\tau$ (scale parameter controlling the degree of orthogonality constraint) in AOP-G. These results show that $\tau$ can be successfully learned from the data and its posterior mean is not extremely small, which is a different feature from that of NeMo imposing nearly orthogonal constraint.
@@
Figure 5 displays the posterior means of the principal functions for each method, where only principal functions having variations, $\int\widehat{f}_{k}(t)^{2}dt$, larger than 0.2, are expressed by a solid line. The number of such principal functions are 10 (NO), 6 (NO-S), 5 (NeMO) and 4 (AOP-G). Also, we present inner produces of the estimated principal functions in Figure 6. From these figures, it is observed that the proposed AOP-G method yields interpretable leading components, effectively capturing dominant temporal variations such as morning and evening population peaks. Notably, the orthogonal priors automatically eliminate unnecessary components by shrinking their amplitudes toward zero, resulting in a sparse low-rank representation that clearly highlights the dominant modes of variation. In contrast, the non-orthogonal approaches (NO and NO-S) suffer from overlapping principal components, even when shrinkage prior is applied. While the NeMO method provides more orthogonal principal functions than NO and NO-S, the first principal function is dominant and the other significant principal functions are relatively similar, compared with those of AOP-G. In fact, the orthogonality measures used in Section 4 are 1.72 (NeMO) and 0.18 (AOP-G), which would indicate that AOP-G provides more orthogonal principal functions.
@@
Figure 4: Trace plot of the posterior samples (left) and posterior and prior densities (right) of the scale parameter $\tau$ to control the strength of orthogonality in AOP-G.
@@
To further investigate the usefulness of the proposed methods, we projected the functional observations into two dimensional principal component score, $(Z_{i1},Z_{i2})$, which are shown in Figure 7. The colors in the scatter plots represent two groups obtained via $k$-means clustering on the original functional data (black and red). We observe that the orthogonal approaches, NeMO and AOP-G result in clearer separation of clusters than the non-orthogonal approaches, NO and NO-S. Further, it can be seen that two clusters are successfully separated only by the first principal component function detected by AOP-G. This would exhibit the advantage of AOG-P to extract meaningful principal functions and provide efficient low-rank representations for functional data.

### v2

@@ Introduction
$f_{1}(t)$ $=\mathbb{I}\Big(0\leq t<\frac{1}{2}\Big)-\mathbb{I}\Big(\frac{1}{2}\leq t\leq 1\Big),\ \ \ \ f_{2}(t)=\sqrt{2}\left[\mathbb{I}\Big(0\leq t<\frac{1}{4}\Big)-\mathbb{I}\Big(\frac{1}{4}\leq t<\frac{1}{2}\Big)\right],$
$\hskip 76.82234ptf_{3}(t)=\sqrt{2}\left[\mathbb{I}\Big(\frac{1}{2}\leq t<\frac{3}{4}\Big)-\mathbb{I}\Big(\frac{3}{4}\leq t\leq 1\Big)\right].$
@@
for the Haar wavelet scenario, where $\mathbb{I}(\cdot)$ is an indicator function. The latent scores for the first three components were generated independently from normal distributions with decreasing variances to reflect dominant leading components: $Z_{i1}\sim{\mathcal{N}}(0,1)$, $Z_{i2}\sim{\mathcal{N}}(0,0.7^{2})$, and $Z_{i3}\sim{\mathcal{N}}(0,0.5^{2})$. The observed functional data $Y_{i}(t)$ were then obtained by adding Gaussian noise with standard deviation $\sigma=1$ to the true signal.
@@
For the generated dataset, we applied the proposed adaptive orthogonal prior with a global inverse-gamma prior on $\tau^{2}$ (we call AO-G), and local inverse-gamma priors on each $\tau_{k}^{2}$ (AO-L), where cubic B-splines with an intercept are used for basis functions. For both priors, we incorporate the sparsity-inducing prior described in Section  2.5. For comparison, we adopted a non-orthogonal prior (NO) which uses a standard normal prior for each coefficients, and a non-orthogonal shrinkage prior (NO-S) applying the horseshoe prior (Carvalho et al., 2009) for coefficients of basis functions to reduce the number of irrelevant principal functions. Moreover, we compared nearly mutually orthogonal (NeMO) processes (Matuk et al., 2022). The (largest) number of possible components was set to $K=10$ in all the methods, and the number of basis functions in NO, NO-S, AO-G and AO-L was set to $L=12$, noting that the results were not sensitive to the choice of the different values for $L$. For NeMO, the threshold value for orthogonal constraint is set to $10^{-4}$, following the default setting of the R package provided by the author’s GitHub repository. Posterior inference was conducted via MCMC, running for 3,000 iterations after discarding 2,000 burn-in iterations.
@@
We first show the result with a single simulation data under scenario (i). In Figure 2, we visualize the estimated principal functions and heatmaps of their inner products, obtained by “NO” (non-orthogonal) and “AO-G” (orthogonal) methods. Overall, the results demonstrated that our proposed orthogonal priors successfully enforced near-orthogonality, effectively selected the number of components, and improved reconstruction accuracy while maintaining reasonable uncertainty quantification. In particular, the orthogonal approaches accurately estimated the true number of principal components (three), with the remaining components being automatically shrunk toward zero. This shrinkage behavior is evident in the estimated principal functions, where only the leading three functions show substantial variation, while the others are effectively eliminated. Furthermore, the heatmaps of inner products provide a clear contrast between the methods: under the proposed orthogonal priors, the off-diagonal elements are nearly zero, indicating that the estimated functions are nearly orthogonal. In contrast, the non-orthogonal methods exhibit large off-diagonal inner products, implying strong correlations among components and reducing interpretability. These findings highlight the advantage of orthogonality-inducing priors in achieving interpretable, parsimonious representations in functional principal component analysis.
@@
    Number of effective components (NC): The number of components with sufficiently large variations, defined as ${\rm NC}=\sum_{k=1}^{K}\mathbb{I}\big\{\int\widehat{f}_{k}^{2}(t)dt>\epsilon\big\}$, where $\widehat{f}_{k}(t)\ (k=1,\ldots,K)$ is the posterior mean of the principal functions. We set $\epsilon=0.1$ in this study.
@@
    Mean squared error (MSE): Estimation accuracy of the mean function $\mu_{i}(t)=Z_{i1}f_{1}(t)+Z_{i2}f_{2}(t)+Z_{i3}f_{3}(t)$, defined as $n^{-1}\int\{\widehat{\mu}_{i}(t)-\mu_{i}(t)\}^{2}dt$, where $\widehat{\mu}_{i}(t)$ is the posterior mean of the mean function.
    Interval score (IS): Coverage performance of the point-wise $95\%$ credible intervals of the mean function $\mu_{i}(t)$, defined in Gneiting and Raftery (2007).
@@
Tables 1 summarizes the averaged values of NC and OG, based on 200 Monte Carlo replications. It shows that the non-orthogonal model without shrinkage (NO) consistently overestimates the number of components, returning the maximum value of $K=10$ in all cases. Although the non-orthogonal model with shrinkage (NO-S) reduces the number of estimated components, it still tends to slightly overestimate when the sample size is large. In contrast, the adaptive orthogonal prior approaches (AO-G and AO-L) provide NC values close to the true number of components, demonstrating the effectiveness of orthogonality-inducing priors in promoting sparsity. On the other hand, NeMO tends to produce NC values larger than the true number of components. Regarding OG, the proposed AO-G and AO-L yield considerably smaller values than the other methods, indicating that the estimated principal functions are nearly orthogonal. On the other hand, both NO and NO-S exhibit large OG values, especially in large sample settings, reflecting strong dependencies among components. Also, it is observed that the OG values of NeMo are slightly larger than those of AO-G and AO-L. These results confirm that the proposed priors not only improve parsimony but also enhance interpretability by ensuring near-orthogonal decomposition.
Table 2 provides the averaged values of MSE and IS, except for NeMO. The reason is that the MSE of NeMO is larger than the other method possibly due to the difference of underlying models (Gaussian process for NeMo and basis functions for the other methods), so that this would not provide fair comparison of the performance. The results show that the proposed AO-G and AO-L priors provides better estimation and inference performance than than non-orthogonal priors, indicating that incorporating adaptive orthogonality enhances not only interpretability but also accuracy of estimation and inference.
@@
Table 1: Average values of estimated number of principal component functions (NC) and orthogonality measures (OG) among principal functions, based on 200 Monte Carlo replications.
Scenario 1 Scenario 2
$n$ NO NOS NeMO AO-G AO-L NO NO-S NeMO AO-G AO-L
50 8.48 3.01 5.08 3.01 3.02 8.35 3.02 5.19 3.00 3.01
NC 100 8.94 3.03 5.43 3.00 3.00 8.65 3.06 5.34 3.00 3.00
200 9.16 3.28 5.34 3.00 3.00 8.94 3.30 5.32 3.00 3.00
50 16.0 2.49 0.84 0.05 0.04 18.1 3.91 0.58 0.05 0.03
OG 100 23.2 3.02 0.87 0.05 0.02 24.3 3.80 0.58 0.05 0.02
200 30.2 4.46 0.85 0.05 0.03 30.5 5.38 0.56 0.04 0.03
Table 2: Average values of mean squared errors (MSE) of the posterior mean of the underlying functions and interval scores (IS) of the $95\%$ point-wise credible intervals of the underlying functions, based on 200 Monte Carlo replications.
Scenario 1 Scenario 2
$n$ NO NOS AO-G AO-L NO NO-S AO-G AO-L
50 0.16 0.13 0.13 0.13 0.16 0.14 0.13 0.13
MSE 100 0.13 0.12 0.11 0.11 0.13 0.12 0.12 0.12
200 0.11 0.11 0.11 0.11 0.12 0.11 0.11 0.11
50 1.85 1.68 1.63 1.62 1.91 1.77 1.70 1.69
IS 100 1.68 1.56 1.54 1.52 1.76 1.66 1.62 1.61
200 1.57 1.49 1.47 1.46 1.65 1.58 1.56 1.55
Sensitivity of the scale parameter for constraint
To check the sensitivity of the scale parameter $\tau^{2}$ controlling the strength of the orthogonality in the proposed AO prior, we evaluated the performance with several choices of $\tau^{2}$. Using the same data generating scenarios, we evaluated the performance by the four evaluation values introduced in the previous section. The results are presented in Table 3. It highlights that simply fixing the scale parameter $\tau^{2}$ at a very small value does not necessarily lead to improved performance. While a small $\tau^{2}$ enforces strong orthogonality in the prior, overly stringent constraints can introduce numerical instability and reduce flexibility, which may deteriorate data fit and result in larger MSE, unnecessarily large number of non-empty functions or inefficient posterior inference (i.e. larger IS). On the other hand, setting $\tau^{2}$ too large weakens orthogonality constraints as expected. As a consequence, the estimated functions may exhibit non-orthogonality as confirmed by larger values of OG. In contrast, the proposed AO-G approach, which estimates $\tau^{2}$ adaptively from the data, is able to automatically learn an appropriate balance between orthogonality and flexibility. The resulting performance demonstrates that adaptive learning of the constraint strength is crucial for achieving both accurate estimation and effective orthogonality, underscoring the necessity of treating $\tau^{2}$ as an inferential quantity rather than a fixed tuning parameter.
Table 3: Performance of the orthogonal prior with fixed $\tau$ and AO-G prior, based on 200 Monte Carlo replications.
$\tau^{2}$
$n$ $10^{-4}$ $10^{-3}$ $0.01$ $0.05$ $0.2$ $0.5$ AO-G
50 5.35 5.17 4.86 3.37 3.00 3.00 3.00
NC 100 6.21 6.26 5.36 3.26 3.00 3.00 3.00
200 7.00 7.05 5.45 3.13 3.00 3.00 3.00
50 0.10 0.11 0.12 0.08 0.05 0.09 0.06
OG 100 0.11 0.10 0.13 0.06 0.04 0.08 0.05
200 0.10 0.10 0.13 0.04 0.04 0.07 0.05
50 0.20 0.21 0.18 0.14 0.13 0.13 0.13
MSE 100 0.22 0.22 0.19 0.12 0.11 0.12 0.11
200 0.23 0.23 0.18 0.11 0.11 0.11 0.11
50 2.02 2.12 1.91 1.67 1.63 1.66 1.63
IS 100 2.09 2.10 1.92 1.57 1.54 1.57 1.54
200 2.16 2.16 1.90 1.49 1.48 1.50 1.48
@@
Here we demonstrate the usefulness of the proposed orthogonal prior in Bayesian FPCA, using the population data collected by NTT Docomo Inc., which is the predominant mobile company in Japan. Following the previous studies (Wakayama et al., 2025; Wakayama and Sugasawa, 2024), we focus on the seven special wards of Tokyo metropolitan area and 500-meter mesh grids covering these areas. While the original data were provided as three-dimensional arrays, representing locations, time points, and dates, we use the aggregated data consisting of average hourly population counts in February 2019 on weekdays, which results in $n=452$ functional observations with $m_{i}\equiv m=24$ discrete observed hours. For $i=1,\ldots,n$, let $Y_{i}^{\ast}(t)$ be the aggregated functional observation, and we use the scaled version $Y_{i}(t)$ defined as $Y_{i}(t)=Y_{i}^{\ast}(t)/\sqrt{m^{-1}\sum_{j=1}^{m}Y_{i}^{\ast}(t_{j})^{2}}$. Figure 3 shows the scaled functional data, which clearly exhibits patterns indicate strong temporal variation, which motivates the use of a low-rank functional representation via FPCA.
@@
We fitted the Bayesian FPCA with four priors, NO, NO-S, NeMO and AO-G, to these functional data. We adopted the same settings for the prior distributions and tuning parameters as used in Section 4, such as $K=10$ (maximum number of principal functions) and $L=12$ (the number of basis functions). We generated 5,000 iterations after 5,000 burn-in periods. In Figure 4, we present the posterior distribution of $\tau$ (scale parameter controlling the degree of orthogonality constraint) in AO-G, and the orthogonality measures for fixed $\tau$. These results show that $\tau$ can be successfully learned from data and its posterior mean is not extremely small, which is a different feature from that of NeMo imposing nearly orthogonal constraint. We observe that the orthogonality measure remains below a stable threshold for values of $\tau$ up to approximately $0.05$, while it increases monotonically for larger values of $\tau$. Notably, the posterior mean of $\tau$ under the proposed method is located around this boundary. Based on the simulation results in Section 4, such choice of $\tau$ would be reasonable to balance orthogonality and efficiency.
@@
Figure 5 displays the posterior means of the principal functions for each method, where only principal functions having variations $\int\widehat{f}_{k}(t)^{2}dt>0.2$ are displayed by solid lines. The number of such principal functions are 10 in NO, 6 in NO-S, 5 in NeMO and 4 in AO-G. Also, we present inner products of the estimated principal functions in Figure 6. From these figures, it is observed that the proposed AO-G method yields interpretable leading components, effectively capturing dominant temporal variations such as morning and evening population peaks. Notably, the orthogonal priors automatically eliminate unnecessary components by shrinking their amplitudes toward zero, resulting in a sparse low-rank representation that clearly highlights the dominant modes of variation. In contrast, the non-orthogonal approaches (NO and NO-S) suffer from overlapping principal components even when shrinkage prior is applied. While the NeMO method provides more orthogonal principal functions than NO and NO-S, the first principal function is dominant and the other significant principal functions are relatively similar, compared with those of AO-G. In fact, the orthogonality measures used in Section 4 are 1.72 (NeMO) and 0.18 (AO-G), which would indicate that AO-G yields more orthogonal principal functions.
@@
Figure 4: Orthogonality measures as a function of scale parameter $\tau$ (left) and posterior and prior densities (right) of the scale parameter $\tau$ to control the strength of orthogonality in AO-G. The dotted line in the left panel indicates the posterior mean of $\tau$.
@@
To further investigate the usefulness of the proposed methods, we projected the functional observations into two dimensional principal component score, $(Z_{i1},Z_{i2})$, which are shown in Figure 7. The colors in the scatter plots represent two groups obtained via two-class $k$-means clustering on the original space (black and red). We observe that the orthogonal approaches, NeMO and AO-G result in clearer separation of clusters than the non-orthogonal approaches, NO and NO-S. Further, it can be seen that two clusters are successfully separated only by the first principal component function detected by AO-G. This would exhibit the advantage of AO-G to extract more meaningful principal functions and provide efficient low-rank representations for functional data.